# An Exposition of Pathfinding Strategies Within Lightning Network Clients

### v1

@@ IV-C2 Channel-Wise Success Probability
Another popular model to estimate $P_{e}$ is based on modeling the relationship between the intended payment amount and the channel capacity [13]. Capacity is the sum of balances across the nodes in the channel. While channel’s capacity is publicly known, the individual balances remain private. A large channel capacity relative to the payment amount intuitively leads to greater probability of success, which is why it can be used within a success probability estimate. Specifically, assuming a uniform distribution for channel liquidity, the failure probability can be expressed as the ratio of payment amount $\mathrm{amt}_{e}$ to the channel capacity $\mathrm{cap}_{e}$, such that the success probability $P_{e}$ can be written as
$P_{e}=1-\frac{\mathrm{amt}_{e}}{\mathrm{cap}_{e}}=\frac{\mathrm{cap}_{e}-\mathrm{amt}_{e}}{\mathrm{cap}_{e}}.$
In this modeling, the success probability $P_{e}$ decreases as the payment amount approaches the channel’s capacity. This model is used in Eclair.
Based on the same assumption, a finer model for $P_{e}$ can be obtained if upper and lower bounds $\mathrm{UB}_{e}$ and $\mathrm{LB}_{e}$ for the available balance (or liquidity) in a channel $e$ are available, as mentioned in [12]. This approach to model the channel success probability is inspired by [14]. If it is known that the (unknown) liquidity balance $x$ satisfies $\mathrm{LB}_{e}\leq x<\mathrm{UB}_{e}$, under a uniform probability assumption, it holds that $P(x=b)=\frac{1}{\mathrm{UB}_{e}-\mathrm{LB}_{e}}$ for any $b$ with $\mathrm{LB}_{e}\leq b<\mathrm{UB}_{e}$, which implies that we can model the probability that a payment of size $\mathrm{amt}_{e}$ through channel $e$ fails as $P(x<\mathrm{amt}_{e})=\sum_{b=\mathrm{LB}_{e}}^{\mathrm{amt}_{e}-1}P(x=b)=\frac{\mathrm{amt}_{e}-\mathrm{LB}_{e}}{\mathrm{UB}_{e}-\mathrm{LB}_{e}}$. Therefore, the success probability can be expressed as
$P_{e}=P(x\geq\mathrm{amt}_{e})=1-P(x<\mathrm{amt}_{e})=\frac{\mathrm{UB}_{e}-\mathrm{amt}_{e}}{\mathrm{UB}_{e}-\mathrm{LB}_{e}}.$
@@
While the previous approaches are based on the assumption of uniform liquidity distribution, assuming a bimodal liquidity distribution is more practical, as the channel liquidity is often unbalanced due to unidirectional flows from the strong sources to strong recipients. This results in liquidity being concentrated either on the local or the remote side [35]. Therefore, the success probability estimation can be based on the assumption of a bimodal distribution [36] for the available channel balance (or channel liquidity) with density function
$P(x)\sim e^{\frac{-x}{\mathrm{s}}}+e^{\frac{x-\mathrm{cap}_{e}}{\mathrm{s}}}.$
Here, $\mathrm{s}$ is a liquidity broadening scale that describes how quickly the bimodal distribution drops off from the extreme balance states of $x=0$ and $x=\mathrm{cap}_{e}$, respectively. A small value of $\mathrm{s}$, relative to channel capacities, represents a strongly bimodal channel balance distribution. In contrast, a large value of $\mathrm{s}$ indicates that the channel balances are more uniformly distributed.
@@ IV-D Attempt Cost/Failure Cost, Hop Cost
These costs are similar to the fee described above. Unlike fees, these costs do not add to the payment amount and are referred to as virtual costs. They consist of two components: fixed (constant) virtual cost and proportional virtual cost.
@@
Pathfinding can also incorporate a virtual cost as one of the weight terms $a_{e}^{i}$ in (2) to optimize the path length. This virtual cost, called hop cost, is used in Eclair.
@@ V-A Shortest Path Algorithms for Additive Costs
Given the Lightning Network graph $G=(V,E)$, many LN clients use Dijkstra’s algorithm Dijkstra$(G,\operatorname{weight}_{a},s)$ (Algorithm 1), to find the cheapest path from source $s\in V$ node (which is, technically, chosen to be the recipient node $r$ of a Lightning payment) to all nodes $u\in V$ in the network (and in particular, to the sender of the LN payment) with respect to an additive total path cost $c(p)$ as in (1), where $\operatorname{weight}_{a}:E\to\operatorname{\mathbb{R}}_{\geq 0}$ is a non-negative channel-specific weight function. On a high level, within the algorithms, distances and predecessors for each vertex $u\in V$ are first initialized. A set $S$ of vertices, whose final shortest-path weights from $s$ have been determined, is maintained. $Q$ is a min-priority queue of vertices, keyed by the total path weight $c_{a}(p)$ as defined in (1). Each vertex $u\in V$ is inserted into the priority queue $Q$, with its priority determined by its initial distance. While there exist vertices in $Q$ to process, the vertex $u$ with the smallest distance estimate $u.c_{a}(p)$ is extracted from $Q$ and the shortest path to $u$ is considered final and added to $S$. For each neighbor $v$ of the vertex $u$, the algorithm attempts to relax the edge $(u,v)$, i.e., the algorithm checks whether the current best-known distance to $v$, $v.c_{a}(p)$, can be improved by traversing via the edge $(u,v)$. If the distance $dist_{a}$ via $u$, $u.c_{a}(p)+weight_{a}(u,v)$, provides a shorter path to $v$, then the predecessor of $v$ can be updated to $u$ and the distance to $v$ can be updated in $Q$, such that $v.c_{a}(p)=u.c_{a}(p)+\operatorname{weight}_{a}(u,v)$.
@@
Timelock $\sum_{e\in p}\operatorname{\Delta^{\text{CLTV}}_{e}}\leq\mathrm{Max\_CLTV\_Limit}$ LND, LDK
Probability $-\sum_{e\in p}\log(P_{e})\leq-\log(Min\_Prob\_Limit)$ LND
Fee $\sum_{e\in p}\mathrm{fee}_{e}\leq\mathrm{Max\_Fee\_Limit}$ LND, LDK
@@
The constraint-adhering version Dijkstra$(G,\operatorname{weight}_{a},y,\mathrm{alpha},s)$ of Algorithm 1 can be interpreted as a greedy algorithm for solving the so-called constrained shortest-path problem (with respect to cost (7) and constraint (8)), a problem class that has been studied by a considerable body of literature [34, Chapter 16], [41, Section 8.4], [42, 43, 44, 45].
While the conventional shortest path problem can be solved, e.g., by Algorithm 1 in a time that is essentially linear in $|E|$ as mentioned above, the constrained shortest-path problem is known to be $\mathcal{NP}$-complete, [34, Appendix B.4],[41, Section 8.4], which implies that, unless $\mathcal{P}=\mathcal{NP}$, no polynomial-time algorithm exists that can solve the constrained shortest-path problem for all problem instances. Thus, the output of Dijkstra$(G,\operatorname{weight}_{a},y,\mathrm{alpha},s)$ will not be able to optimally solve the problem it is designed for solving in general.
@@ V-B Shortest Path Algorithms for Additive+Multiplicative Costs
We discussed in Section IV-C that path cost modelling used in the LND LN client implementation amounts to a additive-plus-multiplicative cost function $c(p)$, which was defined in (5). In order to find paths $p$ that minimize such a cost $c(p)$, LND uses a modified Dijkstra-style algorithm, which we outline in Algorithm 2. In particular, compared to Algorithm 1, the priority queue is keyed based on the additive-plus-multiplicative cost $c(p)$ instead of on a purely additive cost (7), and the algorithm ModDijkstra$(G,\operatorname{weight}_{a},\operatorname{weight}_{m},c_{\mathrm{attempt}},s)$ takes as additional input a multiplicative weight function $\operatorname{weight}_{m}:E\to\operatorname{\mathbb{R}}_{\geq 1}$ with outputs larger than or equal $1$, and an initial cost term factor $c_{\mathrm{attempt}}$.
@@
Algorithm 2 Dijkstra-Style Algorithm Modified for Additive+Multiplicative Cost (5) as used in LND (with colored modification if to impose constraint $\operatorname{x}(p)\leq\mathrm{alpha}$)
@@
         if $(\mathrm{dist}_{a}+\mathrm{dist}_{m})<(v.c_{a}(p)+v.c_{m}(p))$ $\mathrm{cstval}\leq\mathrm{alpha}$ then
@@
              Decrease-Key($Q,v,\mathrm{dist}_{a}+\mathrm{dist}_{m}$)
@@
It turns out that, even without the imposition of any side constraints (8) (no colored modification in Algorithm 2), Algorithm 2 is not guaranteed to find the optimal path cannot always be guaranteed due the multiplicative weight used in $c(p)$ of (5).
@@
However, this is not the optimal path, as the cost of the path $s\to h\to j\to r$ is $4+2+1\cdot 1\cdot 5=11$, which is cheaper than the path returned by Algorithm 2. The algorithm ignores this path during relaxation, as illustrated, due to the nature of the additive-plus-multiplicative cost as in (5), which leads the algorithm to greedily select the predecessor based on the current weight, thereby excluding potentially less expensive routes. Via this counterexample, we conclude that, while the algorithm might often behave reaonsably, the modified Dijkstra-style algorithm ModDijkstra$(G,\operatorname{weight}_{a},\operatorname{weight}_{m},c_{\mathrm{attempt}},s)$ used in LND cannot guarantee to return an optimal path with respect to its modelled cost function (5).
@@ V-C $K$-Shortest Path Algorithms
For this purpose, in particular, Yen’s $K$-shortest path algorithm [46] is used. This algorithm uses a shortest path algorithm (such as Algorithm 1) as a subroutine, but returns a list of $K$ “best’‘ paths with respect to the imposed cost function. If Dijkstra’s algorithm is used as a subroutine, and if no constraints (8) are imposed, the worst-case time complexity of the algorithm is $O(K|V|(|E|+|V|)\log(|V|))$ (for Dijkstra implemented with binary heap) [47, Section 6.3]. This complexity might limit the scalability of the approach in case of further growth of the Lightning Network. Yen’s $K$-shortest path algorithm is currently used only in the Eclair LN client.
@@ VI-A LND
$\mathrm{weight(e)}$ $=\mathrm{fee}_{e}+\mathrm{amt}_{e}\cdot\operatorname{\Delta^{\text{CLTV}}_{e}}\cdot\,\mathrm{riskfactor}$ (9)
@@
In this equation, $\mathrm{fee}_{e}$ and $\operatorname{\Delta^{\text{CLTV}}_{e}}$ are specific to the channel and are obtained through the channel update messages [48] within the network. In LND, the influence of the $\operatorname{\Delta^{\text{CLTV}}_{e}}$ is controlled by scaling it using the $\mathrm{riskfactor}$, which defaults to a value of $15\cdot 10^{-9}$.
$\mathrm{penalty}_{e}=\mathrm{attemptcost}_{e}\cdot\frac{1}{0.5-\frac{0.9\cdot\mathrm{timepref}}{2}}-1.$ (10)
In the above equation, $\mathrm{timepref}$ is the time preference for the payment. It is set to -1 to prioritize optimizing for fees only, set to 1 to prioritize optimizing for reliability only, or set to a value in between for a balanced approach. By default, its value is 0.
@@
The weight function also incorporates the design element attempt cost (Section IV-D). This is a predetermined virtual cost that signifies the cost of a failed payment attempt, denoted as $\mathrm{attemptcost}_{e}$. This serves as a parameter in the weight function to balance the consideration of potentially better routes against the success probability of the payment. The default values of $\mathrm{BaseAttemptCost}$ and $\mathrm{AttemptCostRate}$ are 100 msat and 1000 msat respectively.
Additionally, LND takes into account the success probability of the path $\mathrm{P_{p}}$ as one of its weight function design elements, given by,
$\mathrm{P}_{p}=\prod_{e\in p}P_{e}$ (12)
where $p=(e_{1},\ldots,e_{\operatorname{length}(p)})$ is the path.
It keeps records of past failed payment attempts within a channel, allowing for the tracking of the time elapsed since the last failure. This information is used in calculating success probabilities, underscoring LND’s emphasis on prioritizing reliable routing paths. LND uses two distinct methods for calculating the channel-wise success probability $\mathrm{P_{e}}$ of a channel forwarding a payment: Apriori and Bimodal.
In Apriori, the success probability is influenced by the weight of the last failure. Initially, when a failure occurs, the success probability drops to zero. However, over time, based on the defined $\mathrm{PenaltyHalfLife}$ period, the success probability gradually recovers to its node probability ($\mathrm{nodeProb}_{e}$) via the formula
@@
In Bimodal estimation of success probability, LND assumes a bimodal distribution $\mathrm{P(x)}\sim e^{\frac{\mathrm{-x}}{\mathrm{s}}}+e^{\frac{\mathrm{x-cap_{e}}}{\mathrm{s}}}$ as explained in Section IV-C2. The liquidity broadening scale $\mathrm{s}$ gets the default value of ${3\cdot 10^{8}\,\mathrm{sats}}$.
@@
Upon solving the integral we get the following equation for channel-wise success probability,
@@
where $\mathrm{sa}_{e}$ and $\mathrm{fa}_{e}$ represent the unsettled success and failure amounts, respectively and $\mathrm{cap}_{e}$ is the channel capacity.
From (16), it can be noted that the bimodal estimator provides insights about sending amounts based on previous success and failure amounts, unlike the apriori estimator (13) which does not provide such insights.
The weight function (9) used by LND is founded on the principle of assessing channels with high reliability, characterized by low fees and short timelocks. This rationale is reflected in the weight function’s consideration of fees, timelock, and success probability.
VI-B CLN
Design elements considered in the weight function of CLN are $\mathrm{amount}$, $\mathrm{fee}$, $\mathrm{timelock}$ and $\mathrm{riskfactor}$, leading to the weight function

### v2

@@ IV-C2 Channel-Wise Success Probability
Another popular model to estimate $P_{e}$ is based on modeling the relationship between the intended payment amount and the channel capacity [10, 15]. Capacity is the sum of balances across the nodes in the channel. While channel’s capacity is publicly known, the individual balances remain private. A large channel capacity relative to the payment amount intuitively leads to greater probability of success, which is why it can be used within a success probability estimate. Specifically, assuming a uniform distribution for channel liquidity, the failure probability can be expressed as the ratio of payment amount $\mathrm{amt}_{e}$ to the channel capacity $\mathrm{cap}_{e}$, such that the success probability $P_{e}$ can be written as
$P_{e}=1-\frac{\mathrm{amt}_{e}}{\mathrm{cap}_{e}}=\frac{\mathrm{cap}_{e}-\mathrm{amt}_{e}}{\mathrm{cap}_{e}}.$ (8)
In this modeling, the success probability $P_{e}$ decreases as the payment amount approaches the channel’s capacity. This model is used in the Eclair LN node implementation [15].
Based on the same assumption, a finer model for $P_{e}$ can be obtained if upper and lower bounds $\mathrm{UB}_{e}$ and $\mathrm{LB}_{e}$ for the available balance (or liquidity) in a channel $e$ are available, as mentioned in [13]. This approach to model the channel success probability is inspired by [10]. If it is known that the (unknown) liquidity balance $x$ satisfies $\mathrm{LB}_{e}\leq x<\mathrm{UB}_{e}$, under a uniform probability assumption, it holds that $P(x=b)=\frac{1}{\mathrm{UB}_{e}-\mathrm{LB}_{e}}$ for any $b$ with $\mathrm{LB}_{e}\leq b<\mathrm{UB}_{e}$, which implies that we can model the probability that a payment of size $\mathrm{amt}_{e}$ through channel $e$ fails as $P(x<\mathrm{amt}_{e})=\sum_{b=\mathrm{LB}_{e}}^{\mathrm{amt}_{e}-1}P(x=b)=\frac{\mathrm{amt}_{e}-\mathrm{LB}_{e}}{\mathrm{UB}_{e}-\mathrm{LB}_{e}}$. Therefore, the success probability can be expressed as the conditional probability
$\begin{split}P_{e}&=P(x\geq\mathrm{amt}_{e}\mid\mathrm{LB}_{e}\leq x<\mathrm{UB}_{e})\\ &=1-P(x<\mathrm{amt}_{e}\mid\mathrm{LB}_{e}\leq x<\mathrm{UB}_{e})=\frac{\mathrm{UB}_{e}-\mathrm{amt}_{e}}{\mathrm{UB}_{e}-\mathrm{LB}_{e}}.\end{split}$ (9)
@@
While the previous approaches are based on the assumption of uniform liquidity distribution, assuming a bimodal liquidity distribution is more practical, as the channel liquidity is often unbalanced due to unidirectional flows from the strong sources to strong recipients. This results in liquidity being concentrated either on the local or the remote side [38]. Therefore, the success probability estimation can be based on the assumption of a bimodal distribution [39] for the available channel liquidity with density function proportional to
$P(x)\sim e^{\frac{-x}{\mathrm{s}}}+e^{\frac{x-\mathrm{cap}_{e}}{\mathrm{s}}}$ (10)
within the interval $[0,\mathrm{cap}_{e}]$. Here, $\mathrm{s}$ is a liquidity broadening scale that describes how quickly the bimodal distribution drops off from the extreme balance states of $x=0$ and $x=\mathrm{cap}_{e}$, respectively. A small value of $\mathrm{s}$, relative to channel capacities, represents a strongly bimodal channel balance distribution. In contrast, a large value of $\mathrm{s}$ indicates that the channel balances are more uniformly distributed.
@@
A bimodal distribution for channel liquidity can alternatively be modeled using the quadratic probability density function (PDF) [13] that is proportional to
$P(x)\sim\Big{(}\mathrm{x-\frac{cap_{e}}{2}}\Big{)}^{2}$
within the interval $[0,\mathrm{cap}_{e}]$. The quadratic shape of this PDF has modes at $x=0$ and $x=\mathrm{cap}_{e}$, but decays only quadratically [^3] The documentation [40] erroneously claims that this PDF is exponential, however, it is a quadratic function. towards its minimum at $x=\mathrm{cap}_{e}/2$, whereas the PDF of (10) decays exponentially. Under this assumption, the channel success probability $P_{e}$ is determined as the ratio of the integral of the PDF from the amount to the maximum liquidity, to the integral from the minimum to the maximum liquidity. The $Bimodal$ mode in LDK adopts this approach, and its implementation is detailed in Section VI-C.
@@ IV-D Attempt Cost/Failure Cost, Hop Cost
The notions of attempt, failure or hop costs play similar roles in the edge weight function (2) as the fees described in Section IV-A. Unlike fees, however, these costs are not proportional to the payment amount $\mathrm{amt}_{e}$ and are referred to as virtual costs. They consist of two components: fixed (constant) virtual cost and proportional virtual cost.
@@ V-A Shortest Path Algorithms for Additive Costs
Given the Lightning Network graph $G=(V,E)$, many LN clients use Dijkstra’s algorithm Dijkstra$(G,\operatorname{weight}_{a},s)$ (Algorithm 1), to find the cheapest path from source $s\in V$ node (which is, technically, chosen to be the recipient node $r$ of a Lightning payment) to all nodes $u\in V$ in the network (and in particular, to the sender of the LN payment) with respect to an additive total path cost $c(p)$ as in (1), where $\operatorname{weight}_{a}:E\to\operatorname{\mathbb{R}}_{\geq 0}$ is a non-negative channel-specific weight function. On a high level, within the algorithms, distances and predecessors for each vertex $u\in V$ are first initialized. A set $S$ of vertices, whose final shortest-path weights from $s$ have been determined, is maintained. $Q$ is a min-priority queue of vertices, keyed by the total path weight $c_{a}(p)$ as defined in (1). Each vertex $u\in V$ is inserted into the priority queue $Q$, with its priority determined by its initial distance.
While there exist vertices in $Q$ to process, the vertex $u$ with the smallest distance estimate $u.c_{a}(p)$ is extracted from $Q$ and the shortest path to $u$ is considered final and added to $S$. For each neighbor $v$ of the vertex $u$, the algorithm attempts to relax the edge $(u,v)$, i.e., the algorithm checks whether the current best-known distance to $v$, $v.c_{a}(p)$, can be improved by traversing via the edge $(u,v)$. If the distance $\mathrm{dist}_{a}$ via $u$, $u.c_{a}(p)+weight_{a}(u,v)$, provides a shorter path to $v$, then the predecessor of $v$ can be updated to $u$ and the distance to $v$ can be updated in $Q$, such that $v.c_{a}(p)=u.c_{a}(p)+\operatorname{weight}_{a}(u,v)$.
@@
Timelock $\sum_{e\in p}\operatorname{\Delta^{\text{CLTV}}_{e}}\leq\mathrm{Max\_CLTV\_Limit}$ LND, LDK, Eclair*
Probability $-\sum_{e\in p}\log(P_{e})\leq-\log(Min\_Prob\_Limit)$ LND, LDK
Fee $\sum_{e\in p}\mathrm{fee}_{e}\leq\mathrm{Max\_Fee\_Limit}$ LND, LDK, Eclair*
@@
The constraint-adhering version Dijkstra$(G,\operatorname{weight}_{a},y,\mathrm{alpha},s)$ of Algorithm 1 can be interpreted as a greedy algorithm for solving the so-called constrained shortest-path problem (with respect to cost (11) and constraint (12)), a problem class that has been studied by a considerable body of literature [37, Chapter 16], [45, Section 8.4], [46, 47, 48, 49].
While the conventional shortest path problem can be solved, e.g., by Algorithm 1 in a time that is essentially linear in $|E|$ as mentioned above, the constrained shortest-path problem is known to be $\mathcal{NP}$-complete, [37, Appendix B.4],[45, Section 8.4], which implies that, unless $\mathcal{P}=\mathcal{NP}$, no polynomial-time algorithm exists that can solve the constrained shortest-path problem for all problem instances. Thus, the output of Dijkstra$(G,\operatorname{weight}_{a},y,\mathrm{alpha},s)$ will not be able to optimally solve the problem it is designed for solving in general.
@@ V-B Shortest Path Algorithms for Additive+Multiplicative Costs
We discussed in Section IV-C that path cost modeling used in the LND LN client implementation amounts to an additive-plus-multiplicative cost function $c(p)$, which was defined in (6). In order to find paths $p$ that minimize such a cost $c(p)$, LND uses a modified Dijkstra-style algorithm, which we outline in Algorithm 2. In particular, compared to Algorithm 1, the priority queue is keyed based on the additive-plus-multiplicative cost $c(p)$ instead of on a purely additive cost (11), and the algorithm ModDijkstra$(G,\operatorname{weight}_{a},\operatorname{weight}_{m},c_{\mathrm{attempt}},s)$ takes as additional input a multiplicative weight function $\operatorname{weight}_{m}:E\to\operatorname{\mathbb{R}}_{\geq 1}$ with outputs larger than or equal $1$, and an initial cost term factor $c_{\mathrm{attempt}}$.
@@
Algorithm 2 Dijkstra-Style Algorithm Modified for Additive+Multiplicative Cost (6) as used in LND (with addition of blue lines compared to Algorithm 1 and yellow-background lines if to also impose constraint $\operatorname{x}(p)\leq\mathrm{alpha}$)
@@
         if $(\mathrm{dist}_{a}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+\mathrm{dist}_{m})}<(v.c_{a}(p){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+v.c_{m}(p)})$ $\mathrm{cstval}\leq\mathrm{alpha}$ then
@@
              Decrease-Key($Q,v,\mathrm{dist}_{a}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}+\mathrm{dist}_{m}}$)
@@
It turns out that, even without the imposition of any side constraints (12) (no colored modification in Algorithm 2), Algorithm 2 is not guaranteed to find the optimal path due to the multiplicative weight used in $c(p)$ of (6).
@@
However, this is not the optimal path, as the cost of the path $s\to h\to j\to r$ is $4+2+1\cdot 1\cdot 5=11$, which is cheaper than the path returned by Algorithm 2. The algorithm ignores this path during relaxation, as illustrated, due to the nature of the additive-plus-multiplicative cost as in (6), which leads the algorithm to greedily select the predecessor based on the current weight, thereby excluding potentially less expensive routes. Via this counterexample, we conclude that, while the algorithm might often behave reasonably, the modified Dijkstra-style algorithm ModDijkstra$(G,\operatorname{weight}_{a},\operatorname{weight}_{m},c_{\mathrm{attempt}},s)$ used in LND cannot guarantee to return an optimal path with respect to its modeled cost function (6).
@@ V-C $K$-Shortest Path Algorithms
For this purpose, in particular, Yen’s $K$-shortest path algorithm [50] is used. This algorithm uses a shortest path algorithm (such as Algorithm 1) as a subroutine, but returns a list of $K$ “best’‘ paths with respect to the imposed cost function. If Dijkstra’s algorithm is used as a subroutine, and if no constraints (12) are imposed, the worst-case time complexity of the algorithm is $O(K|V|(|E|+|V|)\log(|V|))$ (for Dijkstra implemented with binary heap) [51, Section 6.3]. This complexity might limit the scalability of the approach in the case of further growth of the Lightning Network. Yen’s $K$-shortest path algorithm is currently used only in the Eclair LN client.
@@ VI-A Lightning Network Daemon (LND)
$\mathrm{weight(e,p)}$ $=\mathrm{fee}_{e}+\mathrm{amt}_{e}\cdot\operatorname{\Delta^{\text{CLTV}}_{e}}\cdot\,\mathrm{riskfactor}$ (13)
@@
where $\mathrm{fee}_{e}$ and $\operatorname{\Delta^{\text{CLTV}}_{e}}$ are specific to the channel and are obtained through the channel update messages [8] within the network, and $\mathrm{amt}_{e}$ is the payment amount through channel $e$. In LND, the influence of the $\operatorname{\Delta^{\text{CLTV}}_{e}}$ is controlled by scaling it using the constant $\mathrm{riskfactor}$, which defaults to a value of $15\cdot 10^{-9}$. In (13), the edge penalty $\mathrm{penalty}_{e}$ is defined as
$\mathrm{penalty}_{e}=\mathrm{attemptcost}_{e}\cdot\frac{1}{0.5-\frac{0.9\cdot\mathrm{timepref}}{2}}-1,$ (14)
where $\mathrm{timepref}$ is a time preference constant for the payment. It is set to $-1$ to prioritize optimizing for fees only, set to $1$ to prioritize optimizing for reliability only, or set to a value in between for a balanced approach. By default, its value is 00. Through (14), LND’s weight function also incorporates the design element attempt cost (Section IV-D) which is defined as
@@
with constants $\mathrm{BaseAttemptCost}$ and $\mathrm{AttemptCostRate}$. This is a predetermined virtual cost that signifies the cost of a failed payment attempt. The default values of $\mathrm{BaseAttemptCost}$ and $\mathrm{AttemptCostRate}$ are $100$ msat and $1000$ msat, respectively.
As can be seen in (13), LND multiplies $\mathrm{penalty}_{e}$ from (14) with the inverse of the success probability $\mathrm{P_{p}}$ of the path $p$, which is a path-dependent cost that prevents the applicability of a standard shortest path algorithm such as Algorithm 1 (cf. Section V-B). LND assumes independence of the channel success probabilities, which leads via $P_{p}=\prod_{e\in p}P_{e}$ (cf. (5)) to an additive-plus-multiplicative total path cost $c(p)$. Therefore, LND uses the modified Dijkstra algorithm as outlined in Algorithm 2 in its pathfinding.
LND keeps records of past failed payment attempts within a channel, allowing for the tracking of the time elapsed since the last failure. This information is used in the estimation of channel success probabilities $P_{e}$, underscoring LND’s emphasis on prioritizing reliable routing paths. LND uses two distinct methods for calculating the channel-wise success probability $\mathrm{P_{e}}$ of a channel forwarding a payment: $Apriori$ and $Bimodal$.
In $Apriori$, the success probability is influenced by the weight of the last failure. Initially, when a failure occurs, the success probability $P_{e}$ drops to 00. However, over time, based on the defined $\mathrm{PenaltyHalfLife}$ period, the success probability gradually recovers to its node probability ($\mathrm{nodeProb}_{e}$) via the formula
@@
In the $Bimodal$ estimation of $P_{e}$, LND assumes a bimodal distribution $\mathrm{P(x)}\sim e^{\frac{\mathrm{-x}}{\mathrm{s}}}+e^{\frac{\mathrm{x-cap_{e}}}{\mathrm{s}}}$ as explained in Section IV-C2. The liquidity broadening scale $\mathrm{s}$ gets the default value of ${3\cdot 10^{5}\,\mathrm{sats}}$.
@@
Upon solving the integral, we obtain the formula
@@
for the channel-wise success probability $P_{e}$, where $\mathrm{sa}_{e}$ and $\mathrm{fa}_{e}$ represent the unsettled success and failure amounts, respectively and $\mathrm{cap}_{e}$ is the channel capacity.
From (19), it can be noted that the $Bimodal$ estimator incorporates information previous success and failure amounts, unlike the $Apriori$ estimator (16), which does not incorporate such information.
As detailed in Table I, LND enforces side constraints during pathfinding in Algorithm 2 to ensure compliance with specific limits. These include $\sum_{e\in p}\operatorname{\Delta^{\text{CLTV}}_{e}}\leq\mathrm{Max\_CLTV\_Limit}$, $\prod_{e\in p}P_{e}\geq 0.01$, and $\sum_{e\in p}\mathrm{fee}_{e}\leq\mathrm{Max\_Fee\_Limit}$. Here, $\mathrm{Max\_CLTV\_Limit}$ and $\mathrm{Max\_Fee\_Limit}$ represent user-defined maximum allowable CLTV and fee limits, respectively.
Overall, the pathfinding strategies used by LND are founded on the principle of assessing channels with high reliability, characterized by low fees and short timelocks. This rationale is also reflected in the weight function’s consideration of fees, timelock, and success probability.
VI-B Core Lightning (CLN)
Design elements considered in the weight function of CLN are success probability, fees and timelock, leading to the weight function